# High-dimensional dynamics in low-dimensional networks

### v1

@@ Discussion
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. Moreover, in the presence of high-dimensional noise or perturbations, low-rank networks often produce high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional responses to high-dimensional perturbations.
The fact that stable, strongly low-rank networks produce low-rank suppression might seem obvious in hindsight: Whenever $W_{0}$ is normal, stability implies that $W$ has a large, negative eigenvalues, so dynamics are highly suppressive in the direction of the corresponding eigenvectors (as in Figures 1 and 2b,c). However, when $W_{0}$ is highly non-normal $W_{0}$, the network still produces low-rank suppression even when eigenvalues are small in magnitude (Supplementary Figure S.2), so this argument alone does not explain low-rank suppression.
The fact that low-rank networks produce high-dimensional responses to high-dimensional inputs might also seem obvious in hindsight: Inputs drive the network, so high-dimensional input should drive high-dimensional activity regardless the internal network structure. However, our condition on the alignment of singular vectors shows that this explanation is not sufficient. When left- and right-singular vectors are not aligned ($\sigma_{P}\ll 1$), high-dimensional inputs drive low-dimensional responses (Figure 2f,g and Supplementary Figure S.2).
Our results have implications for networks in nature and applications. For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 17, 4, 18, 19, 20, 3, 5, 24]. Theoretical studies demonstrating low-dimensional dynamics in low-rank neuronal network models are consistent with some neural recordings showing low-dimensional neural dynamics [41, 42, 43]. However, many of these recordings are made during the presentation of low-dimensional stimuli or motor tasks. A growing number of more recent studies have shown that neural activity is high-dimensional during the presentation of high-dimensional stimuli or tasks [13, 43, 44, 45, 46, 47]. Our results are consistent with the hypothesis that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [13]. Our results are also consistent with observations that neural populations are more sensitive to random perturbations than to perturbations that are aligned with low-dimensional network structure [36].
Beyond neuroscience, our results imply that perturbations or inputs to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space. This result can be used to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
@@ Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equation simulations – except for Figure 5 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 5 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
@@ Acknowledgments
We thank Ashok Litwin-Kumar for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
@@
Figure S.1: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau{\bf{z}}^{\prime}=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure S.2: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 1 except $W_{0}=|c|{\bf{u}}{\bf{v}}^{T}$ where ${\bf{u}}$ and ${\bf{v}}$ are independent random unit vectors. Also, in panel c, we used a perturbation of ${\bf{x}}={\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ as given in Eq. (15)) and in panel e-g, we projected onto ${\bf{v}}$. The network exhibits low-rank suppression in response to perturbations parallel to ${\bf{u}}+(W_{1}{\bf{v}}-{\bf{v}})/|c|$ (panel c) even though the connectivity matrix, $W$, does not have strongly negative eigenvalues (minimum real part of the eigenvalues of $W$ here is $\textrm{re}(\lambda)_{\textrm{min}}=-0.9$ compared to $\textrm{re}(\lambda)_{\textrm{min}}=-10$ in Figure 1). However, the network produces low-dimensional responses to high-dimensional inputs (panel d).
Figure S.3: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (3), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $-0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure S.4: Low-dimensional dynamics in a modular network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the modular network from Figure 3.
@@
Figure S.6: Low-dimensional dynamics in a spatial network. Same as Figure 1e-g except we used the dynamics in Eq. (3) and the spatial network from Figure 4.
Figure S.7: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor ($c=-2$), and the dynamics obey Eq. (3). The instability produced by taking $\rho>1$ generates chaotic dynamics [16, 3].
Figure S.9: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1d–g except ${\bf{x}}(t)$ was made low-dimensional by multiplying it by a matrix with rank-one structure, ${\bf{x}}\leftarrow W_{x}{\bf{x}}$ where $W_{x}$ is a matrix generated identically to, but independently from $W$. The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$.
@@
In this section we review the relationship between our models and analysis and the models and analyses from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 17, 4, 18, 19, 20], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
All three studies quoted above consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model. We next describe the salient properties that distinguish these models from ours.
@@
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (6), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) scale the same in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$.
This difference in scaling is acknowledged by Ostojic et al. [2, 17, 4, 18, 19, 20] who refer to their networks as “weakly low-rank.” Keeping with this terminology, we will use “weakly low-rank” in this section to refer to networks in which the singular values of the low-rank components ($W_{0}$) and high-rank components ($W_{1}$) both scale like $\mathcal{O}(1)$. We additionally use “strongly low-rank” to refer to networks like ours in which the singular values of $W_{0}$ diverge to $\infty$ as $N\to\infty$ (typically, they scale like $\mathcal{O}(\sqrt{N})$) while the singular values of $W_{1}$ scale like $\mathcal{O}(1)$.
To compare weakly and strongly low-rank networks, we can consider the simple rank-one example from Figure 1. Specifically,
$W_{0}=c{\bf{u}}{\bf{u}}^{T}$
while $W_{1}$ has unbiased random entries with variance $\rho/\sqrt{N}$ and dynamics obey Eq. (1). Strongly low-rank networks have $c\gg 1$ while $c\sim\mathcal{O}(1)$ in weakly low-rank networks. In strongly low-rank networks, stability requires that $c<0$, which is critical for low-rank suppression. In weakly low-rank networks, however, it is possible to take $c>0$ without causing instability if $c<1$. When $c>0$, weakly low-rank networks can respond to perturbations aligned to ${\bf{u}}$ more strongly than random perturbations (Supplementary Figure S.11), reversing the trend of low-rank suppression. When $c<0$ in weakly low-rank networks, perturbations in the direction of ${\bf{u}}$ can be suppressed relative to random perturbations, but the effect is weaker than in strongly low-rank networks (Supplementary Figure S.12).

### v2

@@ Discussion
We presented theory and examples showing that effectively low-rank recurrent networks often suppress perturbations aligned with the dominant directions of their connectivity matrices. We also derived conditions under which low-rank networks generate high-dimensional dynamics. We showed that many low-rank structures that arise in nature are consistent with low-rank suppression and high-dimensional dynamics.
Recent, parallel work [10] also demonstrated that low-rank networks can produce suppression and high-dimensional dynamics. In that work, connectivity matrices had a small number of non-zero singular values which were $\mathcal{O}(1)$ in magnitude, in contrast to the large singular values assumed in our models (see Supplementary Text S.1 for a more complete comparison). Other related work showed that trained neural networks can exploit suppression to cancel noise [52].
Low-rank suppression might seem obvious in hindsight: When $W_{0}$ is normal and symmetric, the combination of stability and large singular values implies eigenvalues with large, negative real part, so dynamics are highly suppressive in one direction. Similarly, high-dimensional dynamics might seem like an obvious consequence of high-dimensional inputs. However, the example in Figure 2 coupled with the results in Claims 1 and 2 show that reality is more nuanced: Large negative eigenvalues are not necessary for low-rank suppression, and high-dimensional input is not sufficient for high-dimensional responses.
Our results have implications for networks in nature and applications. For example, in neuroscience, the implications of low-rank recurrent connectivity on neural dynamics is a topic of intense research [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. In neuroscience, our results are consistent with the observation that networks of neurons produce high-dimensional activity in the context of high-dimensional stimuli or tasks, but low-dimensional dynamics in response to low-dimensional stimuli or tasks [16].
Beyond neuroscience, our results imply that perturbations to a low-rank network are more effective when they are delivered non-uniformly across sub-populations or space. More generally, perturbations to a low-rank network are more effective when they are not aligned to the network’s low rank structure. This result can be leveraged to design and test more effective interventions, for example to epidemiological, ecological, or social networks.
@@ Methods
All simulations and analysis were performed in Python using a combination of custom written PyTorch and NumPy code. All differential equations – except for Figure 6 and Supplementary Figure S.10 – were solved using a simple forward Euler scheme with a step size of $dt=0.01$. The simulations in Figure 6 and Supplementary Figure S.10 were solved using a Runge-Kutta scheme adapted from the approach used in previous work [1]. Code to produce all figures is available at
@@ Acknowledgments
We thank Ashok Litwin-Kumar and David G. Clark for helpful conversations and comments on drafts of the manuscript. This work was supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0223 and the National Science Foundation under a Neuronex award number DBI-1707400.
@@
Figure S.1: Response properties of a feedforward network with rank-one structure. Same as Figure 1 except the recurrent network was replace by a feedforward network with dynamics satisfying $\tau d{\bf{z}}/dt=-{\bf{z}}^{\prime}+W{\bf{x}}$. Unlike the recurrent network in Figure 1, the feedforward network is most sensitive to inputs aligned with its low-rank structure, and its dynamics are dominated by one-dimensional variability.
Figure S.2: Dynamics of a network with rank-one structure driven by an input perturbation with one dimensional structure. Same as Figure 1 except the dynamics obeyed $\tau d{\bf{z}}/dt=-{\bf{z}}+W{\bf{z}}+W_{x}\bm{\eta}(t)$ where $\bm{\eta}(t)$ is a realization of the same Gaussian stochastic process used for ${\bf{x}}(t)$ in Figure 1, and $W_{x}$ is an effectively low-rank matrix generated identically to, but independently from $W$. The “effective input”, ${\bf{x}}(t)=W_{x}\bm{\eta}(t)$, is therefore low-dimensional (see green dots in panel a). The network dynamics, ${\bf{z}}(t)$, inherit low-dimensional dynamics from ${\bf{x}}(t)$ (blue dots in panel a), but low-rank suppression and cancellation are still exhibited along the vector, $\bf{u}$}, defining the low-rank structure of $W$ (panels b–d).
Figure S.3: Response properties of a network with independent singular left- and right-singular vectors. Same as Figure 2 except $W_{0}=c{\bf{u}}{\bf{v}}^{T}$ where $\bf{u}$} and $\bf{v}$} are independent random unit vectors. The network behaves similarly to the network in Figure 2 because $\bf{u}$} and $\bf{v}$} are nearly orthogonal.
Figure S.4: Dynamics of an unstable, chaotic network with rank-one structure. Same as Figure 1 except $\rho=2$, the magnitude of $c$ was increased by the same factor (c=2c=-2), and the dynamics obey Eq. (9). The instability produced by taking $\rho>1$ generates chaotic dynamics [27, 3].
Figure S.6: Response properties of a network with biased weights. Same as Figure 1 except we used the dynamics in Eq. (9), and $W_{jk}$ are drawn i.i.d. from a normal distribution with mean $0.5/\sqrt{N}$ and standard deviation $0.5/\sqrt{N}$ where $N=200$.
Figure S.7: Low-dimensional dynamics in a modular network. Same as Figure 1 except we used the dynamics in Eq. (9) and the modular network from Figure 4.
@@
Figure S.9: Low-dimensional dynamics in a spatial network. Same as Figure 1 except we used the dynamics in Eq. (9) and the spatial network from Figure 5.
@@
In this section we review the relationship between our models and the models from previous theoretical work on low-rank networks, specifically work by Ostojic et al. [2, 6, 4, 7, 8, 9], work by Thibeault et al. [1], and work by Landau and Sompolinsky [5, 3]
All of these studies consider recurrent networks with connectivity of the form $W=W_{0}+W_{1}$ where $W_{0}$ is low-rank and $W_{1}$ is a full rank random matrix, just like our model. We next describe the salient properties that distinguish these models from ours.
@@
where $r$ is the rank and each ${\bf{m}}_{\mu}$ and ${\bf{n}}_{\mu}$ are $N\times 1$ vectors with entries that scale like $\mathcal{O}(1)$. Specifically, they are taken to be random vectors and the variance of their entries scales like $\mathcal{O}(1)$. Often, they are taken to be biased random vectors with a non-zero mean that also scales like $\mathcal{O}(1)$. Because of the $1/N$ factor in Eq. (12), the variance of the entries in $W_{0}$ scale like $\mathcal{O}(1/N^{2})$ and, when the entries are biased, the mean entry in $W_{0}$ scales like $\mathcal{O}(1/N)$. Regardless of whether entries are biased, the singular values of $W_{0}$ scale like $\mathcal{O}(1)$ in the models considered by Ostojic and colleagues [2]. Hence, the singular values of the low-rank part ($W_{0}$) and the random part ($W_{1}$) have the same scale in the work of Ostojic and colleagues, in contrast to our models in which the singular values of $W_{0}$ are considered to be asymptotically larger than the singular values of $W_{1}$. This difference in scaling is acknowledged by Ostojic et al. [2, 6, 4, 7, 8, 9] who refer to their networks as “weakly low-rank.”
Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except c=0.5c=-0.5 and $\rho=0.01$.
Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=0.5$ and $\rho=0.01$.
The distinction between strongly and weakly low-rank networks is more nuanced when the full-rank, random part, $W_{1}$, of the connectivity matrix is assumed to be weak or absent. For example, recent work by Mastrogiuseppe et al. [10] considered the case in which the full-rank part is absent ($W_{1}=0$) and the low-rank connectivity matrix, $W=W_{0}$, has $\mathcal{O}(1)$ singular values. While their use of $\mathcal{O}(1)$ singular values is similar to weakly low-rank networks from other work [2], the absence of a random part means that their connectivity matrices are truly low-rank. In that work [10], Mastrogiuseppe et al. showed that the emergence of suppression and high-dimensional dynamics depends on the degree of overlap between the left- and right-singular vectors in ways that are not captured by our asymptotic theory.
For the sake of comparison to the models in [10], let us consider a rank-one network structure like the one in Figure 1. Specifically, we take $W_{1}$ to be random with spectral radius $\rho>0$, and take
$W_{0}=c{\bf{u}}{\bf{u}}^{T}.$
In Figure 1 and its analysis, we assume that $\rho\sim\mathcal{O}(1)$. In this case, a strong (i.e., asymptotically dominating) low-rank structure requires that $|c|\gg 1$. Therefore, stability in this case requires that $c<0$, which produces strong, negative eigenvalues (as in Figure 1; however, consider also the examples in Figure 2 and Supplementary Figure S.3 which are strongly low-rank without large eigenvalues). We could alternatively take the random, full-rank part to be weak: $\rho\ll 1$ (or it could be absent, $\rho=0$, as in [10]). In this case, we can take the low-rank part to be moderate in magnitude ($|c|=\mathcal{O}(1)$) while maintaining an asymptotically dominant low-rank structure because $|c|\gg\rho$ (Supplementary Figure S.11 and S.12).
When $c<0$, the network exhibits a weaker form of low-rank suppression along with high-dimensional responses (Supplementary Figure S.11). Similar examples and results are considered in [10]. In conclusion, some of our theoretical results on low-rank suppression and high-dimensional dynamics require that the low-rank part is large in absolute terms ($|c|\gg 1$), not just relative terms ($|c|\gg\rho$).
Since $|c|=\mathcal{O}(1)$, we can now take $c>0$ while keeping the network stable with a dominant low-rank part (in contrast to strongly low rank networks like Figure 1 where stability requires $c<0$). When $c>0$, inputs aligned to $\bf{u}$} are amplified compared to random perturbations (Supplementary Figure S.12), reversing the trend of low-rank suppression. However, the amplification is weak. In response to high-dimensional input, the network dynamics have a dominant principal component direction, but the dominance is weaker than in other examples we have considered (Supplementary Figure S.12d; compare to Figure 2 and Supplementary Figure S.3).